OpenStudy (anonymous):

Let x = 4. What is the value of y in the equation y = 15 + 11 + x?

OpenStudy (rock_mit182):

that means solve when x = (4); y = 15 + 11 + (4)

OpenStudy (anonymous):

ok

OpenStudy (rock_mit182):

easy :) Doesnt it ?

OpenStudy (rock_mit182):

so y = ?

OpenStudy (anonymous):

i am not sure what the answer is

OpenStudy (rock_mit182):

just add up : 15 + 11 + 4; in order to get the value for y

OpenStudy (rock_mit182):

Im pretty sure you can do it

OpenStudy (anonymous):

=30

OpenStudy (rock_mit182):

;) nice work

OpenStudy (anonymous):

thanks

OpenStudy (rock_mit182):

a piece of cake, no prob ;)

OpenStudy (anonymous):

can you help with another

OpenStudy (anonymous):

Which equation describes this statement? The sum of p and q is equal to 24. Let q = 3. A. p + 3 = 24 B. p – 3 = 24 C. p × 3 = 24 D. p ÷ 3 = 24

OpenStudy (rock_mit182):

first off they say: p + q =24; and q =3 so: p + ? =24

OpenStudy (anonymous):

ok

OpenStudy (rock_mit182):

tell me your answer

OpenStudy (anonymous):

p + 3 = 24

OpenStudy (rock_mit182):

sure is right

OpenStudy (anonymous):

am i right

OpenStudy (rock_mit182):

now if they ask you to find the value of p, is easy as well. take a look: p +3 =24 so: p = 24 -3 which is.. ?

OpenStudy (anonymous):

um idk????

OpenStudy (rock_mit182):

yeah you're right good job

OpenStudy (anonymous):

Let j = 16. What is the value of k in the equation k = 2 × j + 4?

OpenStudy (rock_mit182):

please if you have any other question make a new post ;)

OpenStudy (anonymous):

ok

OpenStudy (rock_mit182):

thanks